Fermi National Laboratory

Accelerator Update for October 6, 2008 - October 8, 2008


Monday Oct 6

The day shift began with the shutdown in progress.

An Antiproton Source (Pbar) expert was given permission at 8:26 AM to access and work in the vault and water cage areas.

Operations restored the Meson CAMAC crates at 8:42 AM.

Operators turned off the Tevatron (TeV) sector A1 LCW pump at 8:53 AM for heat exchanger (HE) work.

A controls expert reported at 10:11 AM that he was beginning work on a VME crate in the Main Injector (MI) enclosure MI-10 that would take down some LCW readbacks.

Operators reported at 12:21 PM that they had locally rebooted the TeV sector B3 BPM VME crate.

The water group reported at 1:18 PM that they were bringing up water column #3 at CUB.

A water group expert reported at 3:34 PM that he was turning on the TeV sector A2 LCW pumps. Later, during the evening shift, the expert turned off all the TeV LCW systems.

Tuesday Oct 7

The midnight shift began with the shutdown in progress.

At 5:21 AM, Operations reported that the Master Sub Station alarmed on FIRUS. Operators informed experts.

A Booster expert reported at 7:06 AM that he had switched the Booster West tunnel sumps and lights to backup power in preparation for the scheduled power outage to that area.

Operations reported that the Booster West Tower power outage began at 7:10 AM.

A Pbar expert was given permission at 8:18 AM to turn on and off pumps and to open and close valves in CUB for the Pbar LCW system.

Operations reported at 8:34 AM that air conditioning technicians were working on the units in TeV sectors D, E, and F service buildings.

At 8:57 AM, Operations reported a leaking hose on a Pbar Debuncher quadrupole magnet (D2Q5). Operators valved out the magnet and contacted a water expert.

A controls expert reported at 10:56 AM that he was rebooting MECAR.

A TeV expert reported at 2:44 PM finding an ice ball on a sector A44 spool piece for a quadrupole magnet (AQ9).

The Booster West power returned at 5:04 PM.

A water expert reported at 6:15 PM that he had completed filling the TeV LCW system. The pumps will remain off for the night.

Cryo system experts reported at 6:48 PM that they had completed their work on wet engines at TeV sectors E3 and F1.

Water experts reported at 7:32 PM that they had completed their work on the Pbar 95 LCW system.

Operators reported at 11:10 PM that they had turned on all of the TeV separators, except for B48V, and put them in conditioning mode.

Wednesday Oct 8

The midnight shift began with the shutdown in progress.

Operations turned off the TeV separators at 7:29 AM.

The Plans for Wednesday

The plans for today are to continue with the shutdown.

Most of the experts report that their machine work is on schedule. A possible problem is with the Recycler’s Pelletron. The Generator Volt Meter (GVM) and its motor have some problems that might require repairs by the company that made the devices.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/08/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory